Find L if P = 26; W = 5
P = 2(L + W)

Answer by ikleyn(52873)

.

Then P/2 = L + W,   or

    26/2 = L + 5

     13  = L + 5

     13 - 5 = L

        8   = L.


ANSWER.  L = 8.

Solved.